# Two-Loop Spacelike Splitting Amplitude for $\mathcal{N}=4$ Super-Yang-Mills Theory

### v1

@@ Abstract
The study of collinear behavior for gauge theories in the spacelike region is of great phenomenological and theoretical importance. We analytically calculate the two-loop spacelike splitting amplitude for the full color $\mathcal{N}=4$ Super-Yang-Mills theory. The result is derived by two complementary methods starting from the known amplitude: one is based on a discontinuity analysis, while the other one is based on analytic continuation. Our result explicitly shows terms that violate naive factorization. However we show that factorization is restored at the level of color-summed unpolarized squared amplitudes at next-to-next-to-next-to leading order. We conjecture that the two-loop tripole terms in the generalized splitting amplitudes in QCD are identical to what we obtain in $\mathcal{N}=4$ super Yang-Mills theory.
@@ I Introduction
Despite these efforts, a conclusive statement about the possibility of factorization breaking or not at the cross-section level from spacelike splitting such as (2) is still missing. In this Letter, we initiate a systematic study in this direction by computing, for the first time, the finite terms in the generalized splitting amplitude in (2) to two loops, as depicted in Fig. 1. We perform the calculation in ${\cal N}=4$ Super Yang-Mills Theory (${\cal N}=4$ SYM), a close cousin of QCD that shares similar infrared behavior in perturbative theory. Building upon the remarkable data for high multiplicities amplitudes in the literature Chicherin:2018old ; Chicherin:2020oor ; Henn:2021cyv ; Henn:2024ngj , we systematically develop techniques for the analytic continuation of high-multiplicity amplitudes from timelike collinear kinematics to spacelike collinear kinematics. We believe such techniques will be useful in future investigations of more complicated factorization-breaking configurations.
@@
This Letter is organized as follows. We first calculate the two-loop five-point massless integrals in the spacelike collinear limit in terms of Goncharov polylogarithms (GPLs) and then assemble the $\mathcal{N}=4$ two-loop five-point amplitude in this limit. These computations provide a solid ground for the colliner factorization violating discussion. In parallel, we analyze the analytic continuation of five-point scattering amplitudes, and explicitly identify the collinear factorization violating terms. This computation agrees with the result from our master integral computations. Based on these computations, we finally derive the generalized two-loop splitting amplitudes, which constitutes our main result.
@@ V Two-loop generalized splitting amplitudes
where $f(\epsilon)=(\psi(1-\epsilon)-\psi(1))/\epsilon$.

### v2

@@ Abstract
The study of collinear behavior for gauge theories in the spacelike region is of great phenomenological and theoretical importance. We analytically calculate the two-loop spacelike splitting amplitude for the full color $\mathcal{N}=4$ Super-Yang-Mills theory. The result is derived by two complementary methods starting from the known amplitude: one is based on a discontinuity analysis, while the other one is based on calculating Feynman integrals analytically in the collinear limit. Our result explicitly shows terms that violate naive factorization. However we show that factorization is restored at the level of color-summed unpolarized squared amplitudes at next-to-next-to-next-to leading order. We conjecture that the two-loop tripole terms in the generalized splitting amplitudes in QCD are identical to what we obtain in $\mathcal{N}=4$ super Yang-Mills theory.
@@ I Introduction
Despite these efforts, a conclusive statement about the possibility of factorization breaking or not at the cross-section level from spacelike splitting such as (2) is still missing. In this Letter, we initiate a systematic study in this direction by computing, for the first time, the finite terms in the generalized splitting amplitude in (2) to two loops, as depicted in Fig. 1. We perform the calculation in ${\cal N}=4$ Super Yang-Mills Theory (${\cal N}=4$ SYM), a close cousin of QCD that shares similar infrared behavior in perturbative theory. Building upon the remarkable data for high multiplicities amplitudes in the literature Chicherin:2018old ; Chicherin:2020oor ; Henn:2021cyv ; Henn:2024ngj ; Agarwal:2023suw ; DeLaurentis:2023nss , we systematically develop techniques for the analytic continuation of high-multiplicity amplitudes from timelike collinear kinematics to spacelike collinear kinematics. We believe such techniques will be useful in future investigations of more complicated factorization-breaking configurations.
@@
This Letter is organized as follows. We first calculate the two-loop five-point massless integrals in the spacelike collinear limit in terms of Goncharov polylogarithms (GPLs) GoncharovGPL and then assemble the $\mathcal{N}=4$ two-loop five-point amplitude in this limit. These computations provide a solid ground for the colliner factorization violating discussion. In parallel, we analyze the analytic continuation of five-point scattering amplitudes, and explicitly identify the collinear factorization violating terms. This computation agrees with the result from our master integral computations. Based on these computations, we finally derive the generalized two-loop splitting amplitudes, which constitutes our main result.
@@ V Two-loop generalized splitting amplitudes
where $f(\epsilon)=(\psi(1-\epsilon)-\psi(1))/\epsilon$ with $\psi$ the digamma function.